(Compound value solving for n) How many years will the following take? a. $520 to grow to $1,348.75 if invested at 10 percent compounded annually b. $33 to grow to $56.70 if invested at 7 percent compounded annually c. $95 to grow to $465.31 if invested at 13 percent compounded annually d. $57 to grow to $69.28 if invested at 5 percent compounded annually a. How many years will it take for $520 to grow to $1,348.75 if invested at 10 percent compounded annually? years (Round to the nearest whole number.) b. How many years will it take for $33 to grow to $56.70 if invested at 7 percent compounded annually? years (Round to the nearest whole number.) c. How many years will it take for $95 to grow to $465.31 if invested at 13 percent compounded annually? years (Round to the nearest whole number) d. How many years will it take for $57 to grow to $69.28 if invested at 5 percent compounded annually? years (Round to the nearest whole number.) Enter your answer in each of the answer boxes.
Answer #1
a Number of years    10 =NPER(10%,0,-520,1348.75)
b Number of years      8 =NPER(7%,0,-33,56.7)
c Number of years    13 =NPER(13%,0,-95,465.31)
d Number of years      4 =NPER(5%,0,-57,69.28)